OpenStudy (anonymous):

How many solutions does the equation have? 3(4x - 2) = 2(6x - 3) A. one solution B. infinite solutions C. no solution

OpenStudy (texaschic101):

3(4x - 2) = 2(6x - 3) --- distribute through the parenthesis 12x - 6 = 12x - 6 if both sides equal, then there is infinite solutions. Now if it would have resulted in x = a number, then there would be 1 solution...and if it would have resulted in 0 = 4 (something incorrect), then there would have been no solutions